OpenStudy (anonymous):

Could I find the two missing angle measures if I know some of the side lengths of a right triangle?

OpenStudy (anonymous):

some means more than 1, right?

OpenStudy (anonymous):

Yeah, I'm pretty sure

OpenStudy (anonymous):

then with sine, cosine, tangent, you can find the angles

OpenStudy (anonymous):

yes, by using SAS congruents and those stuff. like.. a^2 = b^2 + c^2 -2ab Cos A

OpenStudy (anonymous):

im not really sure how to use them to find the angles instead of the other way around

OpenStudy (anonymous):

a^2 = b^2 + c^2 -2ab Cos A

OpenStudy (anonymous):

But if you know it's a right triangle, you don't use the cosine law

OpenStudy (anonymous):

Better only use a² + b² = c²

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

that'll help me find the sides, but i need the angles

OpenStudy (anonymous):

sine theta = opposite/ hypotenuse cosine theta = adjacent / hypotenuse tangent theta = opposite/ adjacent

OpenStudy (anonymous):

Pythagorean theorem does not help to find angles

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

so how do i use the trigonometric ratios to find the angles if i only know the right angle?

OpenStudy (anonymous):

im still confused xD

OpenStudy (anonymous):

then you use inverse sine/cosine/tangent

OpenStudy (anonymous):

okay, how do i use those? im a bit new at this

OpenStudy (anonymous):

The sum of the interior angles of a triangle are equal to 180 degree. To find the third angle of a triangle when the other two angles are known subtract the number of degrees in the other two angles from 180 degree :)

OpenStudy (anonymous):

i understand that, but i only know the 90 degree angle

OpenStudy (anonymous):

Let me show you the example so that you will get it better. :)

OpenStudy (anonymous):

@sweetrascal That has nothing to do with this

OpenStudy (anonymous):

Example: How many degrees are in the third angle of a triangle whose other two angles are 40degree and 65degree ? Answer: 180degree - 40degree -65degree = 75degree

OpenStudy (anonymous):

i only know the 90 degree angle... that's not what im looking for

OpenStudy (anonymous):

\[\sin \theta = \frac{1}{2}\]\[\sin^{-1}\sin \theta = \sin^{-1} \frac{1}{2} \implies \theta = \sin^{-1} \frac{1}{2}\]

OpenStudy (anonymous):

Moneybird will help you further :)

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

so both sides are multiplied by the inverse sine?

OpenStudy (anonymous):

not multiplied on your calculator there is a inverse sine buton

OpenStudy (anonymous):

okay, i get it! thank you!